# Path Evolution Model for Endogenous Channel Digital Twin towards 6G Wireless Networks

### v1

@@ III Applications of Path Evolution Model
  • Multi-domain pilot overhead reduction: With temporal dependencies of path feature in PEM, frequency/time-domain ambiguity [12] for channel acquisition under low pilot densities can be tackled. Thus, pilot overhead in frequency/time-domain can be greatly reduced below channel coherence bandwidth and coherence time limitations. Additionally, beam prediction via path evolution can reduce pilot overhead in spatial-domain [13].
  • Multi-functional sensing: Note that explicit delay and angular information in path features characterize the geometrical relationship among BS, UE and scatters. Instant path features can be used for positioning and the shift of path features can be leveraged for velocity estimation. Meanwhile, with path trajectories maintained in PEM, user tracking can be achieved as well.
High-precision and low-overhead CSI acquisition capability of PEM is investigated by following numerical simulations.
@@ III-A Simulation Setup
In the simulations, precise ray-tracer Wireless Insite is adopted to generate CSI data. As shown in the bottom of Fig. 4, a street scenario is considered, where mobile users are distributed in two environments, namely, env-1 and env-2. Mobility of users is generated by microscopic SUMO simulator [14] with a 20 m/s speed limit, which leads to a maximum Doppler shift of 400 Hz. From the bottom of Fig. 4, it is apparent that environment properties distinctly differ between env-1 and env-2. For communication system configurations, bandwidth is 100 MHz and $16\times 1616 × 16$ Uniform Planar Arrays (UPA) are equipped in BS-1 and BS-2. Signal-to-Noise Ratio (SNR) of channel measurement is set as 10 dB. For the neural networks utilized in PEM and other deep learning based baselines, a training dataset with 1600 samples is collected from env-1. Data augmentation in Sec. II-C is adopted for efficient PEM learning.
@@
III-B PEM Functionality
Firstly, we evaluate the functionality of PEM by comparing the physical and digital path evolutions, which is illustrated in Fig. 5. Here, evolutions of three typical paths in env-1 are taken as examples, including one Line of Sight (LoS) path and two reflection paths. During UE mobility, it can be found that the digital evolutions of the three paths can accurately match their physical counterparts. Moreover, PEM can recognize the death of path-2 and the birth of path-3 when passing from building-2 to building-3. Thus, PEM can provide precise prior knowledge for CSI acquisition in real time. Additionally, the evolution error of path 3 in Fig. 5(c) can be progressively reduced with the updated channel measurements, which exhibits the self-sustainability of PEM.
III-C Time-domain Pilot Overhead Reduction
In this subsection, we aim to leverage PEM to reduce pilot overhead in time-domain by increasing SRS period. With high channel dynamics and large SRS period, continuous time-domain channel prediction is vital to guarantee seamless high spectral efficiency. Hereby, tensor neural Ordinary Differential Equation (ODE) channel prediction network [3] is adopted as a continuous CSI acquisition baseline without CDT. Normalized Mean Square Error (NMSE) of channel prediction schemes under different SRS periods is shown in Fig. 6. It can be found that NMSE of PEM is low and nearly remains unchanged when time-domain pilot overhead is reduced by 5 times. On the contrary, NMSE of the baseline obviously increases with the SRS period. Thus, it is evident that PEM can simultaneously enable low-overhead and high-precision CSI acquisition.
III-D Discussion over Generalization and Endogeneity
Intra-system endogeneity and environment generalization can be achieved in PEM. As depicted in Fig. 6, channel prediction performance of PEM is held in the unseen env-2, which is achieved without any external device. Such environmental generalization capability relies on the generalizability of target path evolution discussed in Sec. II-B and practical data augmentation techniques proposed in Sec. II-C. Additionally, we further consider a Channel Impulse Response (CIR) type CKM [15] as an existing CDT baseline, which achieves the NMSE of -15.37 dB in env-1 and 5.69 dB in env-2. Based on the simulation results, channel acquisition performances of these baselines are obviously degraded when tested in the unseen env-2 due to environment discrepancies, which shows the lack of generalizability.
@@ IV Conclusions and Future Research
CDT is a new paradigm for efficient CSI acquisition in 6G wireless networks. In this paper, PEM is further introduced, which is designed to reflect path evolutions in the physical world. Different from existing vision information-based and CKM-based CDTs, PEM possesses three key advantages of full endogeneity, self-sustainability and environmental generalizability, which are vital to achieving low operation and deployment costs. To facilitate PEM, an intelligent operation framework is proposed, which is light-weighted and compatible with existing massive MIMO systems. The generalization advantages of PEM over other CDTs are rigorously analyzed and efficient PEM learning is proposed to reduce data collection burden practically. Based on extensive numerical simulations, we validate that high-precision and low-overhead CSI acquisition among dynamic environments can be effectively achieved by PEM, which can enable large system capacity for 6G wireless networks.
Although great potential and feasibility have been proved to adopt PEM for efficient CSI acquisition, its effectiveness relies on the specific dynamic nature of the wireless channel and the performance to process the input channel measurements during the operation. Hence, the future research directions span from environment-aware operation design to efficient data processing. Firstly, to optimize the operation performance of PEM, environment-aware update period should be designed by intelligent decision-making. Secondly, multi-band and multi-BS channel measurements are available in Frequency Division Duplexing (FDD) systems and dense networks, which enrich the information for PEM but also bring challenges due to their complicated internal correlations. Then, how to leverage the multi-band and multi-BS channel measurements deserves further investigation. Thirdly, PEM can tackle channel measurement noise with extrinsic information. However, the discrepancies among data modalities bring challenges to multi-modal data fusion. Thus, a multi-modal data fusion framework that fully employs specific physics meanings of extrinsic information requires further investigation.

### v2

@@ III Applications of Path Evolution Model
  • Multi-domain pilot overhead reduction: With temporal dependencies of path feature in PEM, frequency/time-domain ambiguity [10] for channel acquisition under low pilot densities can be tackled. Thus, pilot overhead in the frequency/time-domain can be greatly reduced below channel coherence bandwidth and coherence time limitations. Additionally, beam prediction via path evolution can reduce pilot overhead in spatial-domain [8].
  • Multi-functional sensing: Note that explicit delay and angular information in path features characterize the geometrical relationship among BS, UE, and scatters. Instant path features can be used for positioning and the shift of path features can be leveraged for velocity estimation. Meanwhile, with path trajectories maintained in PEM, user tracking can be achieved as well.
The high-precision and low-overhead CSI acquisition capability of PEM is investigated by the following simulations.
@@ III-A Simulation Setup
In the simulations, the precise ray-tracer Wireless Insite is adopted to generate CSI data. As shown in the bottom of Fig. 4, street scenarios are considered, where mobile users are distributed in env-1 and env-2. Explicitly, env-1 and env-2 exhibit distinct geometric layouts, including building layouts, BS heights, and user positions relative to BS. Additionally, different types of materials are adopted in the buildings in env-1 and env-2. Mobility of users is generated by SUMO simulator [6], where speed limits are set as 20 m/s and 30 m/s in env-1 and env-2, respectively. The communication systems operate at a 6 GHz carrier frequency, where the bandwidth is set as 100 MHz and the number of subcarriers is 138. Uniform Planar Arrays (UPA) with shape $16\times 16$ are both equipped in BS-1 and BS-2. The SNR of channel measurement is set as 10 dB. For the neural network in PEM and deep learning-based baseline, a training dataset with 1600 samples is collected from env-1, which is free of training samples in env-2. Data augmentation in Sec. II-C is adopted during the training of PEM.
@@
III-B PEM Functionality
Firstly, we evaluate the functionality of PEM by comparing the physical and digital path evolutions, which is illustrated in Fig. 5. Here, evolutions of three typical paths in env-1 are taken as examples, including one LoS path and two reflection paths. During UE mobility, it can be found that the digital evolutions of the three paths can accurately match their physical counterparts. Moreover, PEM can recognize the death of path-2 and the birth of path-3 when passing from building-2 to building-3. Thus, PEM can provide precise prior knowledge for CSI acquisition in real time. Additionally, the evolution error of path 3 in Fig. 5(c) can be progressively reduced with the updated channel measurements, which exhibits the self-sustainability of PEM.
III-C Time-domain Pilot Overhead Reduction
In this subsection, we aim to leverage PEM to reduce pilot overhead in time-domain by increasing the SRS period. With high channel dynamics and a large SRS period, continuous time-domain channel prediction is vital to guarantee seamless high spectral efficiency. Hereby, tensor neural Ordinary Differential Equation (ODE) channel prediction network [2] and model-based prediction [9] is adopted as a continuous CSI prediction baseline without CDT, where the length of the historical CSI sequence is 8. Normalized Mean Square Error (NMSE) of channel prediction schemes under different SRS periods is shown in Fig. 6. It can be found that the NMSE of PEM is low and nearly remains unchanged when time-domain pilot overhead is reduced by 5 times, which benefits from the real-time prior knowledge of path feature evolution. Caused by the time-domain ambiguity, the NMSE of the baselines obviously increases when the SRS period is large. Thus, it is evident that PEM can simultaneously enable low-overhead and high-precision CSI acquisition.
III-D Discussion over Generalization and Endogeneity
Intra-system endogeneity and environment generalization can be achieved in PEM. As depicted in Fig. 6, the channel prediction performance of PEM is held in the unseen env-2, which is achieved without any external device. Note that 20% of the user distribution area is blocked in env-2. Thus, the simulation results reveal that the PEM is robust to blockage. Conversely, the channel acquisition performance of the deep learning-based baseline is obviously degraded when tested in the unseen env-2, which shows the lack of generalizability.
@@ IV Conclusions and Future Research
CDT is a new paradigm for efficient CSI acquisition in 6G wireless networks. In this paper, PEM is introduced as a new type of CDT, which is designed to reflect path evolutions in the physical world. Compared to existing vision information-based and CKM-based CDTs, PEM possesses three key advantages of full endogeneity, self-sustainability, and environmental generalizability, which are vital to achieving low operation and deployment costs. To facilitate PEM, an intelligent operation framework is proposed, which is light-weighted and compatible with existing massive MIMO systems. The environment generalizability of PEM is rigorously analyzed, which is realized by progressively addressing the distribution shift. Based on extensive simulations, we validate that high-precision and low-overhead CSI acquisition among dynamic environments can be effectively achieved by PEM, which can enable large system capacity for 6G wireless networks.
Although great potential and feasibility have been proved to adopt PEM for efficient CSI acquisition, its effectiveness relies on the specific dynamic nature of the wireless channel and the performance to process the input channel measurements during the operation. Hence, the future research directions span from environment-aware operation design to efficient data processing. Firstly, to optimize the operation performance of PEM, the environment-aware update period should be designed by intelligent decision-making. Secondly, the accuracy of path extraction in PEM is affected by the measurement resources. With limited measurement resources, how to achieve accurate path extraction in a rich multipath environment needs to be resolved. Thirdly, PEM can tackle channel measurement noise with extrinsic information. However, the discrepancies among data modalities bring challenges to multi-modal data fusion. Thus, a multi-modal data fusion framework that fully employs specific physics meanings of extrinsic information requires further investigation. Fourthly, in complex environments with diffusion paths and moving scatters, the distribution shift of the path features evolution becomes more complicated. Therefore, efficient processing algorithms in PEM based on a more comprehensive distribution shift model merit a deeper exploration.